Do It Best Hardware is offering a 140-piece mechanic's tool set plus a $65 tool chest for $226. What is the cost per tool?
Answer by TimothyLamb(4379)

tool set price = 226 - 65 = 161
---
answer:
price per tool = 161/140 = 1.15
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php